# Robust Deep Reinforcement Learning against Adversarial Behavior Manipulation

### v2

@@ 6 Defense method
The defender’s goal is to prevent the victim from adopting the behavior specified by the adversary. Let $R_{\text{tgt}}\colon\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ be the reward function that is maximized when the victim’s behavior exactly matches the adversary’s specification. We then formulate the defender’s objective as follows:
@@
where $\lambda$ is a hyperparameter that adjusts the trade-off between performance and robustness. $R_{\text{tgt}}$ represents the adversary’s gain, and the second term represents the additional gain achievable by the worst-case adversarial policy. Thus, Equation (10) seeks to minimize this worst-case increase in the adversary’s gain. Since the defender cannot access the adversary’s intended behavior, computing $R_{\mathrm{tgt}}$ is infeasible, and hence Equation (10) cannot be optimized directly.
Time-Discounted Regularization Training (TDRT).
To optimize Equation (10), we prove that the increase in the cumulative reward from the reward function $R_{\text{tgt}}$ in Equation (10) is bounded by the sensitivity of the policy’s action outputs to state changes:
@@
Let $R_{\text{tgt}}\colon\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ be the adversary’s objective reward function, and the discount factor be $\gamma\in(0,1)$. Assume that there exists an upper bound $\bar{R}_{\text{tgt}}\in\mathbb{R}$ for $R_{\text{tgt}}$. Then, the following inequality holds:
@@
Based on Theorem 6.1, we propose a robust training framework, Time-Discounted Regularization Training (TDRT). Let $B=\{(s_{t},t)\}_{t=1}^{N}$ be a mini-batch of state-time pairs, where $t$ is the timestep at which state $s_{t}$ was observed. The defender’s objective is redefined as:
@@ 7 Experiments
We empirically evaluate our proposed method on Meta-World [34](continuous action spaces), MuJoCo [35] (continuous action spaces), and MiniGrid [36] (discrete action spaces, vision-based control). Due to space limitations, this section only reports results for the Meta-World environment, which simulates a robotic arm. Results from the other experiments can be found in Appendix F.
We focus on three key aspects: (i) Attack performance: Can the adversarial policy learned through BIA effectively manipulate the victim’s behavior? (ii) Robustness: Does the policy smoothing in TDRT offer greater robustness against behavior-targeted attacks compared to existing defense methods such as adversarial training? (iii) Original task performance: Can the time-discounting component of TDRT mitigate performance degradation on the original tasks compared to traditional policy smoothing without time discounting?
@@
We consider five opposing task pairs: {window-close, window-open}, {drawer-close, drawer-open}, {faucet-close, faucet-open}, {handle-press-side, handle-pull-side}, and {door-lock, door-unlock}. For example, if the victim originally learned the window-close task, the target policy is defined as a policy that completes the window-open task. Since the reward structures of the Meta-world benchmark are relatively complex, the adversary’s reward function (e.g, reward function of window-open) cannot be obtained by simply negating the victim’s reward function (e.g, reward function of window-close). In this sense, forcing a victim agent trained to perform window-close to execute window-open cannot be achieved through a reward-minimization attack but requires a behavior-targeted attack.
Following the prior works [11, 8], we constrain the set of adversarial states $\mathcal{B}(s)$ using the $L_{\infty}$ norm: $\mathcal{B}(s)\triangleq\left\{\hat{s}\mid\|\hat{s}-s\|_{\infty}\leq\epsilon\right\}$, where $\epsilon$ represents the attack budget. All experiments are performed with $\epsilon=0.3$. See G.2 for the results of changing the attack budgets. States are standardized across all tasks, with standardization coefficients calculated during the training of the victim agent. To learn adversarial policy with BIA, we employed DAC [21] of ILfD and OPOLO [51] for ILfO as IL algorithms, which are variants of GAIL. In BIA, 20 episodes of trajectories generated by the target policy are provided as demonstrations. The target policy is fully trained on the adversary’s target task. For a comparison of attack performance across different target policies, see Appendix G.3.
@@
We compare our proposed attack method with three baselines. (i) Random Attack: This attack perturbs the victim’s state observation by random noise drawn from a uniform distribution. This attack works with no-box access and requires no knowledge about the victim. (ii) Targeted PGD Attack: This naive attack method optimizes falsified states $\hat{s}$ using PGD at each time independently to align the victim’s actions with those of the target policy’s at each time: $\hat{s}=\operatorname*{arg\,min}_{\hat{s}}d(\pi(\cdot|\hat{s}),\pi_{\text{tgt}}(\cdot|s))$. PGD requires white-box access to the victim’s policy, giving the adversary an advantage not available in our proposed method. The detailed explanation and analysis of Targeted PGD are provided in Appendix I.4.2. (iii) Target Reward Maximization Attack: This attack leverages Lemma B.1 to learn an adversarial policy that maximizes the cumulative reward obtained by the victim from the adversary’s reward function $R_{\text{adv}}$: $\nu^{*}=\operatorname*{arg\,max}_{{\pi\circ\nu}}^{M}[R_{\text{adv}}(s,a,s^{\prime})]$. We used two methods for training the adversarial policy: SA-RL[11] (black-box attack) and PA-AD[8] (white-box attack). This attack requires access to the adversary’s reward function. Thus, this attack also gives the adversary an advantage not available in our proposed method.
@@
Defense methods against behavior-targeted attacks have not yet been proposed. Therefore, as baselines, we employ defense methods against untargeted attacks: ATLA-PPO[11], PA-ATLA-PPO[8], RAD-PPO[32], WocaR-PPO[24], and SA-PPO[7]. ATLA-PPO and PA-ATLA-PPO are methods that utilize adversarial training. RAD-PPO is a regret-based defense method that learns policies to minimize regret, defined as the difference between the rewards under non-attack and attack conditions. WocaR-PPO is a defense method that learns policies to maximize the worst-case rewards and applies regularization to the smoothness of policies, specifically in critical states where rewards significantly decrease. SA-PPO aims to increase the smoothness of the policy’s action outputs by a regularizer. The difference between TDRT-PPO and SA-PPO is that SA-PPO does not apply time discounting in the regularization. For detailed explanations of the baselines, see Appendix I.2.
Table 1: Comparison of attack performances. Each value represents the average episode reward $\pm$ standard deviation over 50 episodes. Each parenthesis indicates (Access model, Adversary’s knowledge). Under the limited-knowledge setting, BIA’s attack performance is competitive with that of baseline methods that assume an greater adversary knowledge.
@@
We present the results in Table 2. The attack rewards represent the cumulative reward of the adversary’s task obtained by the victim under attacks. The target rewards are the cumulative reward obtained by the target policy, which is directly trained with the adversary’s reward function, serving as the upper bound for attack rewards.
BIA achieves an attack performance comparable to more advantaged attacks, such as Target Reward Maximization (Rew Max/SA-RL, PA-AD), which has access to the adversary’s reward function. This demonstrates that BIA can effectively attack using demonstrations alone, without requiring any reward modeling of the adversary’s task. The attack performance of BIA-ILfO is nearly identical to that of BIA-ILfD. We attribute this to the deterministic nature of state transitions in the MetaWorld tasks: without access to the victim’s actions, the transitions can be sufficiently predicted. While we used 20 episodes as demonstrations for BIA, results in Appendix G.1 confirm that the attack performance remains nearly unchanged even when the demonstrations are reduced to four episodes.
@@
In all tasks, TDRT-PPO and SA-PPO achieve the greatest robustness against attacks, indicating that policy smoothing is effective against behavior-targeted attacks. For the drawer-close task, SA-PPO attains exceptionally high robustness, albeit with a substantial drop in original task performance (see Table 3). Additional experiments in Appendix H.2 examine how varying the smoothing coefficient trades off robustness and original task performance.
Adversarial training methods, including ATLA-PPO and PA-ATLA-PPO, are still vulnerable to behavior-targeted attacks in most tasks. As noted by [52, 53], adversarial training is effective only against the reward-minimization attack anticipated during training. Thus, it remains vulnerable to behavior-targeted attacks. Although the regret-based approach RAD-PPO demonstrates relatively higher robustness in some tasks than adversarial training, it still relies on reward-based metrics and, therefore, does not offer sufficient protection against behavior-targeted attacks. WocaR-PPO, which applies smoothing to a subset of states, achieves moderate robustness but underperforms SA-PPO and TDRT-PPO, both of which uniformly smooth across the entire state space. Further experiments on robust training efficiency are presented in Appendix H.3.
@@
Best Attack Reward ()($\downarrow$)( ↓ )
@@
Clean Reward ()($\uparrow$)( ↑ )
@@
TDRT-PPO achieves higher clean rewards than SA-PPO in all tasks. Together with the results in Table 2, this demonstrates that while uniform regularization across the entire trajectory in SA-PPO sacrifices original performance, time discounting in TDRT-PPO mitigates performance degradation while preserving robustness.
@@ 8 Conclusion and Limitations
This work introduces the Behavior Imitation Attack (BIA), which manipulates victim behavior through perturbed state observations under limited access to the victim’s policy. We also introduced Time-Discounted Regularization Training (TDRT), the first defense method specifically designed for behavior-targeted attacks, which achieved robustness without compromising original performance.
While our work contributes valuable insights, it also has limitations. Adversarial policy is known to be less effective in high-dimensional state spaces, such as image inputs [8] (see Appendix F.2). Efforts to overcome this require white-box access[8] and remain unresolved. Moreover, although TDRT empirically exhibits robustness against behavior-targeted attacks, it lacks certified guarantees. In scenarios where higher reliability requirements are imposed, TDRT may prove insufficient.
@@ A.1 Attack Methods
In reward-minimization attacks, the adversary’s goal is to minimize the cumulative reward received by the victim. The most basic approaches are gradient-based methods. Huang et al. [6] proposed an attack method that uses the Fast Gradient Signed Method (FGSM) [54] to compute adversarial perturbations that prevent the victim from choosing optimal actions. Pattanail et al. [9] performed a more powerful attack by creating adversarial perturbations to minimize the value estimated by the Q-function. Gleave et al. [10] proposed an attack that assumes a two-agent environment. This method creates an adversarial agent that severely degrades the victim’s performance. Sun et al. [37] point out that previous works lack stealth and proposed the Critical Point Attack and Antagonist Attack, which achieve effective attacks within very few steps. Qiaoben et al. [38] classify existing adversarial attacks against RL agents in the function space and propose an attack method based on a two-stage optimization derived from the theoretical analyses. Sun et al. [37] and Qiaoben et al. [38] are similar to ours in that they attempt to bring the victim’s policy closer to a specific policy. However, these objectives differ from ours, and they are implemented via a PGD attack that requires white-box access. Duan et al. [55] proposed an attack that targets the distribution of the victim’s policy. However, their method is a reward-minimization attack and differs from ours. Specifically, they intended to induce large deviations from the original victim’s distribution and indirectly reduce the victim’s cumulative reward. Because their goal is not to shift the policy toward any specific target distribution, their attack differs from the behavior-targeted attack.
@@
The enchanting attack aims to lure the victim agent to reach a predetermined target state. Lin et al. [40] first proposed this type of attack. In their approach, the adversary generates a sequence of states and actions that cause the victim to reach the target state. The adversary then crafts a sequence of perturbations to make the victim perform the sequence of actions. Tretschk et al. [41] proposed an enchanting attack where adversarial perturbations to maximize adversarial rewards are heuristically designed for the attack purpose, thereby leading the victim to the specified target state. Buddareddygari et al. [56] proposed a different enchanting attack using visual patterns placed on physical objects in the environment so that the victim agent is directed to the target state. Unlike [40] and [41], which perturb the victim’s state observations, this attack alters the environmental dynamics. Ying et al. [42] proposed an enchanting attack with a universal adversarial perturbation. When any state observations are modified with this perturbation, the victim agent is forced to be guided to the target state. All of these enchanting attacks require white-box access to the victim’s policy.
@@
The behavior-targeted attack aims to manipulate not only the final destination but also the victim’s behavior in a more detailed. Hussenot et al. [16] proposed a behavior-targeted attack that forces the victim to select the same actions as the policy specified by the adversary. More specifically, this attack precomputes a universal perturbation for each action so that the victim who observes the perturbed states takes the same action as the adversarially specified policy. One limitation of this attack is that the computational cost of precomputing such universal perturbations can be high when the action space is large or continuous. Since the cost required for pre-computing perturbations is significant, applying this attack is challenging. Boloor et al. [17] proposed a heuristic attack specifically designed for autonomous vehicles. They formulated an objective function with a detailed knowledge of the target task, which is not applicable to behavior-targeted attacks for general tasks. Bai et al. [19, 18] investigated an attack to manipulate the victim’s behavior that follows the adversary’s preference for the behavior. This attack method can be applied to any environment. However, one limitation of this method is that it usually requires thousands of labeled preference state-action sequences to specify the behavior that the adversary requests to follow. Therefore, this attack is not practical. We remark that all existing behavior-targeted attacks also require white-box access to the victim’s policy.
@@
Several studies have proposed poisoning attacks against RL agents, which intervene during the victim’s training phase. Sun et al. [43] proposed VA2C-P, a poisoning attack framework that can adapt without access to environment dynamics. They proposed both untargeted attacks and targeted attacks that aim to align the victim’s policy with a target policy. Rakhsha et al. [57] theoretically investigated the attack problem in both offline planning and online RL settings with tabular MDPs. Rangi et al. [44] also provided theoretical insights into the fundamental limits of poisoning attacks in episodic reinforcement learning. Xu et al. [45, 46] proposed more practical black-box attacks that do not require knowledge of environment dynamics or the victim’s learning algorithm. As we focus on attacking trained victims, our threat model differs from those considered in these studies.

### v3

@@ 6 Defense method
In this section, we propose a countermeasure against behavior-targeted attacks. In this attack, the adversary’s objective is defined independently of the victim’s reward. As a result, the victim can be steered toward unintended and potentially malicious behaviors even when the victim’s cumulative reward does not decrease. Therefore, existing defenses that aim to prevent reward degradation may fail to provide sufficient robustness under this threat. To address this limitation, we incorporate the adversary’s gain as a safety metric.
@@
Let $R_{\text{tgt}}\colon\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ be the reward function that is maximized when the victim’s behavior exactly matches the adversary’s specification. We then formulate the defender’s objective as follows:
@@
where $\lambda$ is a hyperparameter that adjusts the trade-off between performance and robustness. $R_{\text{tgt}}$ represents the adversary’s gain, and the second term represents the additional gain achievable by the worst-case adversarial policy. Thus, equation 10 seeks to minimize this worst-case increase in the adversary’s gain. We note that equation 10 subsumes prior defense objectives against reward-minimization attacks via the choice of $R_{\mathrm{tgt}}$. In particular, setting $R_{\mathrm{tgt}}=-R$ makes the second term a regularizer that suppresses reward drops under attack, yielding a standard robustness objective.
Since the defender cannot access the adversary’s intended behavior, equation 10 cannot be optimized directly. In particular, the adversary’s objective is specified independently of the victim’s reward, so the defender cannot compute the target reward $R_{\mathrm{tgt}}$. Consequently, the gradients required to optimize equation 10 are unavailable.
Time-Discounted Robust Training (TDRT).
To optimize equation 10, we prove that the increase in the cumulative reward from the reward function $R_{\text{tgt}}$ in equation 10 is bounded by the sensitivity of the policy’s action outputs to state changes:
@@
Let $R_{\text{tgt}}\colon\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ be the adversary’s objective reward function, and the discount factor be $\gamma\in(0,1)$. Assume that there exists an upper bound $\bar{R}_{\text{tgt}}\in\mathbb{R}$ for $R_{\text{tgt}}$. Then:
@@
Based on Theorem 6.1, we propose a robust training framework, Time-Discounted Robust Training (TDRT). Let $B=\{(s_{t},t)\}_{t=1}^{N}$ be a mini-batch of state-time pairs, where $t$ is the timestep at which state $s_{t}$ was observed. The defender’s objective is redefined as:
@@ 7 Experiments
We empirically evaluate our proposed method on Meta-World (Yu et al., 2020) (continuous action spaces), MuJoCo (Todorov et al., 2012) (continuous action spaces), and MiniGrid (Chevalier-Boisvert et al., 2018) (discrete action spaces, vision-based control). Due to space limitations, this section only reports results for the Meta-World environment. Other experiments can be found in Appendix F.
We focus on three key aspects: (i) Attack performance: Can the adversarial policy learned through BIA effectively manipulate the victim’s behavior? (ii) Robustness: Does the policy smoothing in TDRT offer greater robustness against behavior‑targeted attacks compared to existing defense methods such as adversarial training? (iii) Original task performance: Can the time‑discounting component of TDRT mitigate performance degradation on the original tasks compared to traditional policy smoothing without time discounting?
@@
We consider five opposing task pairs: {window-close, window-open}, {drawer-close, drawer-open}, {faucet-close, faucet-open}, {handle-press-side, handle-pull-side}, and {door-lock, door-unlock}. For example, if the victim originally learned the window-close task, the target policy is defined as a policy that completes the window-open task. Since the reward structures of the Meta-World benchmark are relatively complex, the adversary’s reward function (e.g., reward function of window-open) cannot be obtained by simply negating the victim’s reward function (e.g., reward function of window-close). In this sense, forcing a victim agent trained to perform window-close to execute window-open cannot be achieved through a reward-minimization attack but requires a behavior-targeted attack.
Following the prior works (Zhang et al., 2021; Sun et al., 2022), we constrain the set of adversarial states $\mathcal{B}(s)$ using the $L_{\infty}$ norm: $\mathcal{B}(s)\triangleq\left\{\hat{s}\mid\|\hat{s}-s\|_{\infty}\leq\epsilon\right\}$, where $\epsilon$ represents the attack budget. All experiments are performed with $\epsilon=0.3$. See G.2 for the results of changing the attack budgets. States are standardized across all tasks, with standardization coefficients calculated during the training of the victim agent. To learn adversarial policies with BIA, we employed DAC (Kostrikov et al., 2018) of ILfD and OPOLO (Zhu et al., 2020) for ILfO as IL algorithms, which are variants of GAIL. In BIA, the demonstrations consist of 20 episodes of trajectories generated by the target policy, which is fully trained on the adversary’s target task. We also vary the number of demonstrations in Appendix G.1. The results show that performance changes little even when the number of episodes is reduced to four. Appendix G.3 also provides a comparison of attack performance across different target policies.
@@
We compare our proposed attack method with three baselines. (i) Random Attack: This attack perturbs the victim’s state observation by random noise drawn from a uniform distribution. This attack works with no-box access and requires no knowledge about the victim. (ii) Targeted PGD Attack: This naive attack method optimizes falsified states $\hat{s}$ using PGD at each time independently to align the victim’s actions with those of the target policy’s at each time: $\hat{s}=\operatorname*{arg\,min}_{\hat{s}}d(\pi(\cdot|\hat{s}),\pi_{\text{tgt}}(\cdot|s))$. PGD requires white-box access to the victim’s policy, giving the adversary an advantage not available in our proposed method. The detailed explanation and analysis of Targeted PGD are provided in Appendix I.4.2. (iii) Target Reward Maximization Attack: This attack leverages Lemma B.1 to learn an adversarial policy that maximizes the cumulative reward obtained by the victim from the adversary’s reward function $R_{\text{adv}}$: $\nu^{*}=\operatorname*{arg\,max}_{{\pi\circ\nu}}^{M}[R_{\text{adv}}(s,a,s^{\prime})]$. We used two methods for training the adversarial policy: SA-RL(Zhang et al., 2021) (black-box attack) and PA-AD(Sun et al., 2022) (white-box attack). These attacks require access to the adversary’s reward function. Thus, they give the adversary an advantage not available in our proposed method.
@@
Defense methods against behavior-targeted attacks have not yet been proposed. Therefore, as baselines, we employ defense methods against untargeted attacks: ATLA-PPO(Zhang et al., 2021), PA-ATLA-PPO(Sun et al., 2022), RAD-PPO(Belaire et al., 2024), WocaR-PPO(Liang et al., 2022), and SA-PPO(Zhang et al., 2020b). ATLA-PPO and PA-ATLA-PPO are methods that utilize adversarial training. RAD-PPO is a regret-based defense method that learns policies to minimize regret, defined as the difference between the rewards under non-attack and attack conditions. WocaR-PPO is a defense method that learns policies to maximize the worst-case rewards and applies regularization to the smoothness of policies, specifically in critical states where rewards significantly decrease. SA-PPO aims to increase the smoothness of the policy’s action outputs by a regularizer. The difference between TDRT-PPO and SA-PPO is that SA-PPO does not apply time discounting in the regularization. For detailed explanations of the baselines, see Appendix I.2.
Table 1: Comparison of attack performances. Each value represents the average episode reward $\pm$ standard deviation over 50 episodes. Each parenthesis indicates (Access model, Adversary’s knowledge). Under the limited-knowledge setting, BIA’s attack performance is competitive with that of baseline methods that assume greater adversary knowledge.
@@
We present the results in Table 1. The attack rewards represent the cumulative reward of the adversary’s task obtained by the victim under attacks. The target rewards are the cumulative reward obtained by the target policy, which is directly trained with the adversary’s reward function, serving as the upper bound for attack rewards. We also report attack success rates in Table 10 by using the task-success criterion in Meta-World. These results exhibit the same trend.
BIA achieves an attack performance comparable to more advantaged attacks, such as Target Reward Maximization (Rew Max/SA-RL, PA-AD), which has access to the adversary’s reward function. This demonstrates that BIA can effectively attack using demonstrations alone, without requiring any reward modeling of the adversary’s task. The attack performance of BIA-ILfO is nearly identical to that of BIA-ILfD. We attribute this to the deterministic nature of state transitions in the Meta-World tasks: without access to the victim’s actions, the transitions can be sufficiently predicted. While we used 20 episodes as demonstrations for BIA, results in Appendix G.1 confirm that the attack performance remains nearly unchanged even when the demonstrations are reduced to four episodes.
@@
In all tasks, TDRT-PPO and SA-PPO achieve superior robustness against attacks, indicating that policy smoothing is effective against behavior-targeted attacks. For the drawer-close task, SA-PPO attains exceptionally high robustness, albeit with a substantial drop in original task performance (see Table 3). Additional experiments in Appendix H.2 examine how varying the smoothing coefficient trades off robustness and original task performance.
Adversarial training methods, including ATLA-PPO and PA-ATLA-PPO, are still vulnerable to behavior-targeted attacks in most tasks. As noted by (Korkmaz, 2021; 2023), adversarial training is effective only against the reward-minimization attack anticipated during training. Thus, it remains vulnerable to behavior-targeted attacks. Although the regret-based approach RAD-PPO demonstrates relatively higher robustness in some tasks than adversarial training, it still relies on reward-based metrics and, therefore, does not offer sufficient protection against behavior-targeted attacks. WocaR-PPO, which applies smoothing to a subset of states, achieves moderate robustness but underperforms SA-PPO and TDRT-PPO, both of which uniformly smooth across the entire state space. Further experiments on robust training efficiency are presented in Appendix H.3.
@@
Best Attack Reward $(\downarrow)$
@@
Clean Reward $(\uparrow)$
@@
TDRT-PPO achieves higher clean rewards than SA-PPO in all tasks. Together with the results in Table 2, this demonstrates that while uniform regularization across the entire trajectory in SA-PPO sacrifices original performance, time discounting in TDRT-PPO mitigates performance degradation while preserving robustness. In Appendix H.2, we further study the effect of the regularization coefficient. Across a wide range of coefficients, TDRT-PPO consistently achieves higher clean rewards than SA-PPO.
@@ 8 Conclusion and Limitations
This work introduces the Behavior Imitation Attack (BIA), which manipulates victim behavior through perturbed state observations under limited access to the victim’s policy. We also introduced Time-Discounted Robust Training (TDRT), the first defense method specifically designed for behavior-targeted attacks, which achieved robustness without compromising original performance.
Limitation.
While our work contributes valuable insights, it also has limitations. Adversarial policy is known to be less effective in high-dimensional state spaces, such as image inputs (Sun et al., 2022) (see Appendix F.2). Efforts to overcome this require white-box access (Sun et al., 2022) and remain unresolved. Moreover, although TDRT empirically exhibits robustness against behavior-targeted attacks, it lacks certified guarantees. In scenarios where higher reliability requirements are imposed, TDRT may prove insufficient.
9 Acknowledgements
This work is partially supported by JST JPMJNX25C2, JPMJKP24C3, JPMJCR23M4, JPMJCR21D3, JSPS 23H00483, and 120251002. We gratefully acknowledge the insightful comments and suggestions provided by the anonymous reviewers.
10 Ethics statements
This paper focused on adversarial attacks on reinforcement learning and their defense methods, aiming to improve the reliability of deep reinforcement learning. Our contribution lies in proposing attack methods and defenses that are envisioned for real-world scenarios. However, our proposed methods still face challenges in practical real-world applications. Our research fully complies with legal and ethical standards, and there are no conflicts of interest. Throughout this study, we utilized only publicly available benchmarks. No private datasets were used in this research.
We have provided detailed algorithms for our proposed methods. The benchmarks used in our experiments are public and accessible to everyone. As described above, we have made substantial efforts to ensure the reproducibility of our results.
@@ A.1 Attack Methods
In reward-minimization attacks, the adversary’s goal is to minimize the cumulative reward received by the victim. The most basic approaches are gradient-based methods. Huang et al. (2017) proposed an attack method that uses the Fast Gradient Signed Method (FGSM) (Goodfellow et al., 2014) to compute adversarial perturbations that prevent the victim from choosing optimal actions. Pattanaik et al. (2018) performed a more powerful attack by creating adversarial perturbations to minimize the value estimated by the Q-function. Gleave et al. (2020) proposed an attack that assumes a two-agent environment. This method creates an adversarial agent that severely degrades the victim’s performance. Sun et al. (2020) point out that previous works lack stealth and proposed the Critical Point Attack and Antagonist Attack, which achieve effective attacks within very few steps. Qiaoben et al. (2024) classify existing adversarial attacks against RL agents in the function space and propose an attack method based on a two-stage optimization derived from the theoretical analyses. Sun et al. (2020) and Qiaoben et al. (2024) are similar to ours in that they attempt to bring the victim’s policy closer to a specific policy. However, these objectives differ from ours, and they are implemented via a PGD attack that requires white-box access.Duan et al. (2025) proposed an attack that targets the distribution of the victim’s policy. However, their method is a reward-minimization attack and differs from ours. Specifically, they intended to induce large deviations from the original victim’s distribution and indirectly reduce the victim’s cumulative reward. Because their goal is not to shift the policy toward any specific target distribution, their attack differs from the behavior-targeted attack.
@@
The enchanting attack aims to lure the victim agent to reach a predetermined target state. Lin et al. (2017) first proposed this type of attack. In their approach, the adversary generates a sequence of states and actions that cause the victim to reach the target state. The adversary then crafts a sequence of perturbations to make the victim perform the sequence of actions. Tretschk et al. (2018) proposed an enchanting attack where adversarial perturbations to maximize adversarial rewards are heuristically designed for the attack purpose, thereby leading the victim to the specified target state. Buddareddygari et al. (2022) proposed a different enchanting attack using visual patterns placed on physical objects in the environment so that the victim agent is directed to the target state. Unlike (Lin et al., 2017) and (Tretschk et al., 2018), which perturb the victim’s state observations, this attack alters the environmental dynamics. Ying et al. (2023) proposed an enchanting attack with a universal adversarial perturbation. When any state observations are modified with this perturbation, the victim agent is forced to be guided to the target state. All of these enchanting attacks require white-box access to the victim’s policy.
@@
The behavior-targeted attack aims to manipulate not only the final destination but also the victim’s behavior in a more detailed manner. Hussenot et al. (2020) proposed a behavior-targeted attack that forces the victim to select the same actions as the policy specified by the adversary. More specifically, this attack precomputes a universal perturbation for each action so that the victim who observes the perturbed states takes the same action as the adversarially specified policy. One limitation of this attack is that the computational cost of precomputing such universal perturbations can be high when the action space is large or continuous. Since the cost required for pre-computing perturbations is significant, applying this attack is challenging. Boloor et al. (2020) proposed a heuristic attack specifically designed for autonomous vehicles. They formulated an objective function with a detailed knowledge of the target task, which is not applicable to behavior-targeted attacks for general tasks. Bai et al. (2025; 2024) investigated an attack to manipulate the victim’s behavior that follows the adversary’s preference for the behavior. This attack method can be applied to any environment. However, one limitation of this method is that it usually requires thousands of labeled preference state-action sequences to specify the behavior that the adversary requests to follow. Therefore, this attack is not practical. We remark that all existing behavior-targeted attacks also require white-box access to the victim’s policy.
@@
Several studies have proposed poisoning attacks against RL agents, which intervene during the victim’s training phase. Kiourti et al. (2020) introduced TrojDRL, a framework for evaluating backdoor attacks in DRL, where the adversary intercepts and manipulates states during training to steer the agent toward a predefined target policy. Sun et al. (2021) proposed VA2C-P that adapts without access to environment dynamics and supports both untargeted attacks and targeted attacks that align the victim’s policy with a target policy. Rakhsha et al. (2020) theoretically analyzed poisoning attacks in both offline planning and online RL with tabular MDPs, while Rangi et al. (2022) further clarified the fundamental limits of poisoning attacks in episodic reinforcement learning. Other works (Xu et al., 2023; Xu and Singh, 2023; Li et al., 2024b) developed more practical black-box poisoning attacks that do not require knowledge of the environment dynamics or the victim’s learning algorithm. More recently, Cui et al. (2024) proposed BadRL, a sparse poisoning attack that dynamically generates sample-specific triggers and injects them only at high attack-value states. All of these methods aim to steer the victim’s policy toward an attacker-specified target policy. However, our threat model explicitly prohibits the adversary from accessing or intervening in the victim’s training process, whereas poisoning attacks inherently rely on such access. Consequently, these attack settings fall outside the scope of our problem formulation.